OpenStudy (anonymous):

complete the square, type answer as an ordered pair: x^2-10x-4=0

OpenStudy (anonymous):

x^2-10x-4+29-29

OpenStudy (anonymous):

??

OpenStudy (anonymous):

x^2-10x-4+29-29 = (x-5)^2 -29 =0

OpenStudy (anonymous):